Mark is three times as old as Luke. In 5 years Mark will be two times as old as Luke is in 5 years. How old is each one now?

Svetlana is four times as old as Boris. In 10 years Svetlana will be three times as old as Boris is in 10 years. How old is each one now?

Masato is twice as old as Kim. Ten years ago, Masato was three times as old as Kim was ten years ago. How old is each one now?

Answer by ankor@dixie-net.com(22740)

Mark is three times as old as Luke. In 5 years Mark will be two times as old as Luke is in 5 years.
How old is each one now?
:
let m = Mark's present age
Let L = Luke's
Write an equation for each statement
:
"Mark is three times as old as Luke."
m = 3L
:
"In 5 years Mark will be two times as old as Luke is in 5 years."
m + 5 = 2(L+5)
m + 5 = 2L + 10
m = 2L + 10 - 5
m = 2L + 5
From the 1st statement, replace m with 3L
3L = 2L + 5
3L - 2L = 5
L = 5 yrs is Luke's present age
then
m = 3(5)
m = 15 yrs is Mark's present age
:
Confirm this in the statement
""In 5 years Mark will be two times as old as Luke is in 5 years."
15 + 5 = 2(5 + 5);
:
:
You can do these other two problems in the same way,
:
"Svetlana is four times as old as Boris."
s = 4b
:
"In 10 years Svetlana will be three times as old as Boris is in 10 years."
s + 10 = 3(b+10)
How old is each one now?
:
:
See if you can do this one completely on your own, difficulty? ankor@att.net
:
Masato is twice as old as Kim. Ten years ago, Masato was three times as old as Kim was ten years ago.
How old is each one now?